The financial statements of P&G are presented in Appendix B. The Company’s complete annual report, including the notes to the financial statements, is available at P&G Annual Reports. Refer to P&G’s financial statements and the accompanying notes to address the following:

How does P&G value its inventories? Which inventory costing method does P&G use as a basis for reporting its inventories?


How does P&G report its inventories in the balance sheet? In the notes to its financial statements, what three descriptions are used to classify its inventories?


What costs does P&G include in Inventory and Cost of Products Sold?


What was P&G’s inventory turnover in 2014? What is its gross profit percentage? Evaluate P&G’s inventory turnover and its gross profit percentage.


Please click on the link provided in the question. Im needing help
Answer #1

Valuation of Inventory:

Inventories are valued at the lower of cost or market value.

Methods of Valuation of Inventories:

Inventories are valued at FIFO, LIFO and Average -Cost Method.

  • Product-related inventories are maintained on the first-in, first-out method.
  • Minor amounts of product inventories, including certain cosmetics and commodities, are maintained on the last-in, first-out method.
  • The cost of spare part inventories is maintained using the average-cost method.

Reporting of Inventories in Balance Sheet:

P&G Reports its Inventories in the Balance Sheet as per the Valuation Method prescribed in its notes mentioned above.

Following three descriptions are used by P&G to classify the inventories in Financial Statements:

INVENTORIES
Materials and supplies 1,742
Work in process 684
Finished goods 4,333
Total inventories 6,759

Cost included by P&G in its Inventories and Cost of Product Sold:

Materials & Supplies comprises purchase cost of direct material and supplies. Work in progress comprises proportionate cost of direct material and supplies along with proportionate manufacturing labor, direct overheads expenses. Finished Goods contains the all the cost which are directly related to manufacture finished goods excluding selling and administrative cost.

Cost of products sold is primarily comprised of direct materials and supplies consumed in the manufacture of product, as well as manufacturing labor, depreciation expense and direct overhead expense necessary to acquire and convert the purchased materials and supplies into finished product. Cost of products sold also includes the cost to distribute products to customers, inbound freight costs, internal transfer costs, warehousing costs and other shipping and handling activity.
When we include selling and administrative cost in cost of product sold, we get cost of sales of particular finished product.

Inventory Turnover, GP Ratio and its evaluation (2014) :

Inventory Turnover = Net Sales / Average Inventory

Average Inventory= (Ending Inventory + Beginning Inventory) / 2

Average Inventory = (6759+6909)/2 = $6834

Inventory Turnover = 83062/6834 = 12.15 Times

GP Percentage: 48.9% (Given)

Evaluation:
Inventory Turnover=>
P&G are earning 12.15 times of our inventory maintained averagely throughout the year.
Gross Profit Percentage=> P&G's Operating Income is 48.9% of its Net Sales.